At a certain college, it is estimated that approximately 19% of the students ride bicycles to school. Would you consider this to be valid estimate if, in a random sample of 85 college students, 20 are found to ride bicycles to class

1
Expert's answer
2022-06-07T11:22:04-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.19"

"H_a:p\not=0.19"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c =1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{\dfrac{20}{85}-0.19}{\sqrt{\dfrac{0.19(1-0.19)}{85}}}\approx1.0645"

Since it is observed that "|z|=1.0645< 1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z>1.0645)=0.287102," and since "p=0.287102>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than "0.19," at the "\alpha = 0.05" significance level.